• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name adaq

    Log entry time 07:27:40 on April 17, 2011

    Entry number 355886

    keyword=Beam Time Accounting
    Hall A

    04/16/11 1600-2400
    Experiment #: e08-009
    Shift Leader: S. Sirca
    Shift Workers: A. Shahinyan (TO), M. Mihovilovic (3rd)

    16001700180019002000210022002300Total Time
    Acceptable Beam
    in Use
    0.950.930.190.000.000.000.000.002.07
    Beam Available
    but Not in Use
    0.020.020.590.330.000.000.000.000.96
    Beam Not Available
    or Unacceptable
    0.030.050.210.671.001.001.001.004.96
    Accelerator Con-
    figuration Change
    0.000.000.000.000.000.000.000.000.00
    Experiment
    Ready
    1.001.001.001.001.001.001.001.008.00
    Planned Config-
    uration Change
    0.000.000.000.000.000.000.000.000.00
    Unplanned
    Experiment Down
    0.000.000.000.000.000.000.000.000.00
    Hall not
    ready reasons
    Target0.000.000.000.000.000.000.000.000.00
    Data Acq0.000.000.000.000.000.000.000.000.00
    Other0.000.000.000.000.000.000.000.000.00
    HRSL0.000.000.000.000.000.000.000.000.00
    HRSR0.000.000.000.000.000.000.000.000.00
    HALL_OFF0.000.000.000.000.000.000.000.000.00
    Bigbite0.000.000.000.000.000.000.000.000.00

    Comments: Reason for experiment not ready: power glitch at 18:50 that knocked out the HRS magnets and the cryo target

    Data Entry by: Simon Sirca